OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 2x + 5 A. (5, 10) and (2, 9) B. (5, 15) and (2, 9) C. (5, 15) and (9, 2) D. (15, 5) and (2, 9)

OpenStudy (anonymous):

Please Help ;3;

OpenStudy (anonymous):

i recommend just plugging in the values and seeing if both sides of the equation are equal.

OpenStudy (anonymous):

so your apost to multiply Oh and wait i have a file to attach

OpenStudy (anonymous):

nvm i dont have a file

OpenStudy (anonymous):

plug in an x and y value from one of the choices into the equation

OpenStudy (anonymous):

i'll show you A as an example

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

now try that for choice B and see if the values match

OpenStudy (anonymous):

y is 5 and x is 15 so 5=2(15 5=20?

OpenStudy (anonymous):

actually that's x is 5 and y is 15. The first number in an ordered pair is always the x value and y is the second value

OpenStudy (anonymous):

Oh,

OpenStudy (anonymous):

so 15=2(5)

OpenStudy (anonymous):

don't forget the +5

OpenStudy (anonymous):

it's 15=2(5)+5

OpenStudy (anonymous):

so 15+2 is 17 and +5+5?

OpenStudy (anonymous):

27?

OpenStudy (anonymous):